# Logos as a Well-Tempered Pre-train for Sign Language Recognition

### v1

@@ 3.1 Dataset Characteristics
The Logos dataset contains 199,668 videos, divided into 80.7% in the train and 19.3% in the test sets. Videos have a resolution of at least 720 pixels on the minimum side at a 30 FPS frame rate. About 62% of videos are in FullHD format. The total duration of the dataset video is 221.4 hours, with 104.7 hours representing the demonstration of signs themselves and the rest being fragments before and after the sign demonstration. The dataset contains 2,863 unique gloss classes combined into 2,004 grouped VSSign classes with 35 to 737 samples per class. The dataset was recorded by 381 signers of various age categories, 41% of them are 30-40 years old, and 88% are female (Figure 2g,h). We do not limit crowdsourcers by age and gender, and such an uneven distribution reflects the demographics of signers who wish to participate in the project. All signers passed an exam confirming their Russian Sign Language proficiency.
The Logos dataset includes the Slovo public dataset Kapitanov et al. (2023) with the renewed annotations. The pipeline for collecting new data repeats the one used for the Slovo dataset, except for the extended gloss selection stage, the new VSSigns grouping stage, and the train-test split stage.
@@ 3.3 VSSigns Grouping
First, we trained a baseline model on the dataset with ungrouped glosses, and processed 2,863 sign template videos with the model. Using confidence of prediction classes for the templates videos, we identified the 10 most similar templates for each one. Deaf experts compared each of the identified videos to their templates and labeled matching ones as VSSigns.
@@ 3.4 Train-test Split
We aim to maintain an 80/20 ratio for the train and test data split applied to both the number of signers and the number of samples for each sign. Given that the number of signs recorded by different signers differs, the dataset split confirming all these requirements hardly has a strict resolution. We applied a dynamic programming algorithm to find the best approximation.
@@ 4.1 Datasets
In addition to the extensive Logos dataset, we selected two widely used ISLR datasets as examples of low-resource datasets: the Turkish Sign Language (TSL) dataset AUTSL Sincan and Keles (2020) and the American Sign Language (ASL) dataset WLASL Li et al. (2020). Their key characteristics are shown in Table 1. The WLASL dataset has a large number of glosses but is relatively small with 21,083 samples, averaging about 10 samples per class. The AUTSL dataset contains more samples but has a limited vocabulary and number of signers.
@@ 4.2 Sign Language Recognition Pipeline
Our experimental setup is based on  Kvanchiani et al. (2024). The authors explore various training aspects to propose the optimal ISLR pipeline. They use MViTv2-S Li et al. (2022) as a backbone, a fully connected (FC) layer for classification, a cross-entropy classification loss with label smoothing, and sign timeline boundary regression as an auxiliary task. The backbone was initialized with Kinetics-400 pre-train. The pipeline processes $32\times 224\times 224$ frame chains, randomly sampled from the input video with a step of 2 frames. We implement an auxiliary boundary regression task as follows. The sign’s ground truth boundary timestamps are rescaled relative to the sampled clip: the clip length is set as 1, the clip start is set as 0 for the sign start point, and the clip end is set as 0 for the sign endpoint. Alongside the classification heads, we add an extra FC layer with two output channels for the sign start and end regression. Its output and scaled ground truth values are mapped to $(-1,1)$ using the formula $y=2\sigma(x)-1$, where $\sigma(x)$ is the sigmoid function, to diminish the influence of sign boundaries that are outside the clip. We use mean squared error loss to train this regression head. The total loss function is calculated as a weighted sum of the classification and regression losses: $L=L_{cls}+2.5L_{regr}$. We evaluate the model using a top-1 instance-based accuracy metric: the ratio of the correctly classified samples to the total samples number.
@@ 4.3 Multi-dataset Co-training Method
Different national sign languages contain different signs for different words. Each dataset has its own label space with no common taxonomy. Therefore, they cannot be directly mixed for simultaneous training.
@@
At the training stage, we use CutMix Yun et al. (2019) and Mixup Zhang (2017) inter-sample regularization strategies. They can not be applied to the mixed batch because labels of different languages cannot be mixed. We use the same language-specific gate to split the mixed batch into language-specific sub-batches before applying these augmentations and then merge the resulting samples back into one batch.
@@ 5.1 Transfer learning experiments
Transfer learning:
@@
Multi-dataset co-training:
@@
Full dataset (baseline) 95.25 62.44
@@
The presented extensive Logos dataset was used as a pre-train for transfer learning tasks. The AUTSL and WLASL datasets were taken as examples of low-resource datasets. First, we trained the separate baseline models on the Logos, AUTSL, and WLASL datasets using the same setup (Section 4). Then, we examined the applicability of the Logos pre-trained model for transfer learning to smaller AUTSL and WLASL datasets. With the model backbone initialized from the Logos pre-train, we evaluated two transfer learning strategies: (a) training all model weights and (b) freezing the pre-trained encoder and training only the classification head. The Logos pre-train substantially improves the model accuracy compared to training from scratch (Table 2).
@@
These experiments demonstrate the possibility of transfer learning from the extensive Logos dataset to other sign languages with only a limited amount of training data. Below, we show that a large Logos dataset size is critical for the encoder quality (Section 5.4) and that VSSigns label grouping improves it (Section 5.5).
5.2 Cross-lingual Multi-dataset Co-training Results
@@
A single model, produced by a multi-dataset co-training, far surpasses the accuracy on low-resource datasets of the models, separately trained on each dataset from scratch, and also surpasses individual models trained using conventional transfer learning (Table 2). Moreover, results for the WLASL dataset are far above existing SOTA metrics444according to https://paperswithcode.com/ and other papers referring to the datasets in question, see Table 4. As for the AUTSL dataset, note that not only models with accuracy better than ours Ryumin et al. (2023); Jiang et al. (2021), but other leading models use ensembling, pose recognition, depth map (or some of the above). In contrast, our model uses a single stream that takes only RGB input. It confirms that the co-training method is promising for low-resource sign languages.
@@
BSL-1K Albanie et al. (2020) 46.9
SignBERT Hu et al. (2021a) 54.7
SAM-SLR Jiang et al. (2021) 98.5 58.7
One Model is Not Enough Hrúz et al. (2022) 96.4
BEST Zhao et al. (2023) 54.6
SignBERT+ Hu et al. (2023) 55.6
NLA-SLR Zuo et al. (2023) 61.3
SL-GDN Miah et al. (2023) 96.5
ST-GCN Papadimitriou and Potamianos (2023) 96.7
Audio-visual … Ryumin et al. (2023) 98.6
HWGAT Patra et al. (2024) 95.8 48.5
StepNet Shen et al. (2024a) 61.2
MViTv2 2024 (our baseline) 96.58 60.88
Multi-dataset 2025 (ours) 97.81 66.82
@@ 5.3 The Encoder Generalization Ability Check
We examined the hypothesis that an encoder pre-trained on the Logos dataset does not produce universal sign features but can only recognize the signs from the train set. When applied to another language, the model maps these signs to the most similar target language signs, as in the approach of Wei and Chen (2023). To emulate this hypothesis, we processed the WLASL train set with the Logos pre-trained model and built the map by associating the assigned Logos labels with the most frequent WLASL ground truth labels. Then, we applied the same model to the WLASL test set and substituted the resulting Logos labels with WLASL labels using the map instead of training a target language classification head. We repeated the same experiment with the AUTSL dataset.
@@
5.4 The Importance of the Dataset Size
pre-train Top-1 accuracy
@@ 5.4 The Importance of the Dataset Size
WLASL 93.16 67.9
Table 6: The importance of the pre-train dataset size for cross-language transfer learning. Results for both whole and truncated versions of the AUTSL and WLASL datasets using pre-training on the Logos dataset and more low-resource WLASL and AUTSL ones.
@@ 5.5 The Effect of VSSigns Grouping
Train Test Top-1 accuracy
Whole non-VSSigns VSSigns
grouped grouped 97.90 97.49 98.33
not grouped grouped 97.44 97.10 97.79
not grouped not grouped 87.02 97.10 76.51
@@
Logos pre-train Top-1 accuracy
@@
We investigated the contribution of our approach with grouping labels of visually similar signs in obtaining a high-quality encoder. We trained the classifier on the Logos dataset, using unique pairs of ungrouped and grouped labels as classes. It formed 2,880 ungrouped gloss classes instead of 2,004 grouped VSSign classes in the baseline Logos annotation. Each ungrouped label has a unique associated grouped label, so the model, trained on the ungrouped labels, can be evaluated on grouped labels.
@@
Table 7 shows that such a model has worse quality than the baseline model, trained on grouped VSSign labels. Notably, the degradation is observed even for signs that are not VSSigns and have unique grouped labels. Furthermore, Table 8 shows that VSSigns grouping results in more effective transfer learning to other sign languages.
6 Limitations
This work is limited by MviT baseline architecture and ISLR cross-language transfer learning as a downstream task. Additional research on other large-scale pre-train datasets, low-resource target datasets, and other downstream tasks, including continuous SL, is needed to generalize the conclusions. We consider it as a perspective.
The Logos dataset has a biased age and gender distribution due to the demographics of the project participants and race distribution due to the natural bias of RSL native users. It can limit its applicability.
7 Conclusions
@@
The paper examines two aspects of the isolated sign language recognition (ISLR) task: cross-language SL model training, including transfer learning, and approaches to handling visually similar signs (VSSigns). To explore these issues, this work presents Logos, a new publicly available Russian Sign Language dataset, the most extensive ISLR dataset by the number of signers and one of the largest available datasets while also the largest RSL dataset in size and vocabulary. It is shown that a model, pre-trained on the Logos dataset can be used as a universal encoder for other language SLR tasks, including few-shot learning. The cross-language transfer learning methods are evaluated, and it is demonstrated that the method of multi-dataset co-training with multiple language-specific classification heads improves SL models for low-resource datasets the most, compared to the conventional “pre-train and finetune” method. The key feature of the Logos dataset is the explicit annotation of visually similar sign groups. With its use, we show that explicitly grouping VSSign labels benefits trained model quality as a video encoder for downstream tasks, such as transfer learning to other sign languages. Based on the proposed contributions, we outperform current state-of-the-art results for the WLASL dataset and get competitive results for the AUTSL dataset, with a single stream model processing solely RGB video.
@@
All crowdworkers provided informed consent, authorizing the processing and publication of the collected data. To save contributors’ privacy, we use anonymized user hash IDs. We do not restrict the participation of signers under 18, provided parental consent was obtained during the registration, in compliance with the Civil Code of the Russian Federation [^5] https://ihl-databases.icrc.org/en/national-practice/federal-law-no-152-fz-personal-data-2006. Compensation for completed tasks was aligned with the average salary of a sign language interpreter proportionate to the time invested. We have verified that the Slovo dataset, incorporated into Logos, adheres to these ethical standards. The dataset is made available exclusively for research purposes. Nonetheless, we acknowledge the potential misuse, such as identifying individuals or enabling large-scale surveillance.

### v2

@@ 3.1 Dataset Characteristics
The Logos dataset contains 199,668 videos recorded by 381 signers (deaf individuals, professional interpreters, sign language teachers, and CODA/SODA [^2] Child of Deaf Adult, Sibling of Deaf Adult/Deaf person). The total duration of the dataset video is 221.4 hours, with 104.7 hours representing the demonstration of signs themselves and the rest being fragments before and after the sign demonstration. The dataset contains signs for 2,863 of the most commonly used lemmas in the Russian general vocabulary, combined into 2,004 grouped VSSign classes with 35 to 737 samples per class.
The Logos dataset includes the Slovo public dataset kapitanov2023SlovoRussianSign with the renewed annotations, amended with VSSign classes.
More details on the dataset’s characteristics are provided in Appendix A.1.
@@ 3.2 Gloss Selection
3.3 Data Collection
The Logos data collection pipeline includes signer selection, video collection, video validation, and sign time interval annotation stages that coincide with the ones of the Slovokapitanov2023SlovoRussianSign pipeline. See Appendix A.2 for details.
@@ 3.4 VSSigns Grouping
First, we trained a baseline model on the dataset with ungrouped glosses and processed 2,863 sign template videos with the model. Using confidence of prediction classes for the template videos, we identified the 10 most similar templates for each one. Deaf experts compared each template video with its 10 counterparts and marked pairs that differ only in non-manual components. Sign pairs matched by the majority of 5 experts were annotated as VSSigns.
@@ 3.5 Train-test Split
We aim to maintain an 80/20 ratio for the train and test data split applied to both the number of signers and the number of samples for each sign. Given that the number of signs recorded by different signers differs, the dataset split confirming all these requirements hardly has a strict resolution. We applied a dynamic programming algorithm to find the best approximation. See Appendix A.3.
@@ 4.1 Datasets
In addition to the proposed large-scale Logos dataset, we selected two widely used ISLR benchmarks as target datasets for transfer learning: the Turkish Sign Language (TİD) dataset AUTSL sincan2020AUTSLLargeScale and the American Sign Language (ASL) dataset WLASL li2020WLASL. The WLASL dataset offers a large vocabulary but is relatively small in size, comprising an average of approximately 10 samples per class. The AUTSL dataset provides more samples but includes a more limited vocabulary and fewer signers. We also include MM-WLAuslan shen2024MMWLAuslanMultiViewMultiModal, one of the most extensive publicly available ISLR datasets, as an additional pre-training baseline to compare against Logos. Key characteristics of these datasets are summarized in Table 1.
@@ 4.2 Sign Language Recognition Pipeline
Our experimental setup is based on  kvanchiani2024TrainingStrategiesIsolated. The authors explore various training aspects to propose the optimal ISLR pipeline. They use MViTv2-S li2022mvitv2 as a backbone, a fully connected (FC) layer for classification, a cross-entropy classification loss with label smoothing, and sign timeline boundary regression as an auxiliary task. The backbone was initialized with Kinetics-400 pre-train. The pipeline processes $32\times 224\times 224$ frame chains, randomly sampled from the input video with a step of 2 frames. We implement an auxiliary boundary regression task as follows. The sign’s ground truth boundary timestamps are rescaled relative to the sampled clip: the clip length is set as 1, the clip start is set as 0 for the sign start point, and the clip end is set as 0 for the sign endpoint. Alongside the classification heads, we add an extra FC layer with two output channels for the sign start and end regression. Its output and scaled ground truth values are mapped to (1,1)(-1,1) using the formula $y=2\sigma(x)-1$, where $\sigma(x)$ is the sigmoid function, to diminish the influence of sign boundaries that are outside the clip. We use mean squared error loss to train this regression head. The total loss function is calculated as a weighted sum of the classification and regression losses: $L=L_{cls}+2.5L_{regr}$. We evaluate the model using a top-1 instance-based accuracy metric: the ratio of the correctly classified samples to the total samples number.
@@ 4.3 Multi-dataset Co-training Method
Different national sign language datasets have their own label spaces with no common taxonomy. As a result, they cannot be directly combined for joint training.
@@
At the training stage, we use CutMix yun2019CutMixRegularizationStrategy and Mixup zhang2018MixupEmpiricalRisk inter-sample regularization strategies. They can not be applied to the mixed batch because labels of different languages cannot be mixed. We use the same language-specific gate to split the mixed batch into language-specific sub-batches before applying these augmentations and then merge the resulting samples back into one batch.
@@ 5.1 Transfer learning experiments
Transfer learning from Logos dataset:
@@
Multi-dataset co-training with Logos dataset:
@@
Full dataset (baseline) 97.25 62.44
@@
The presented extensive Logos dataset was used as a pre-train for transfer learning tasks. We used relatively small AUTSL and WLASL datasets as modeling examples of low-resource datasets for transfer learning. These datasets were also selected because benchmark results are available for comparison. Additionally, we created reduced versions of each dataset, limited to 10, 3, and 1 sample per class, for more challenging low-data experiments.
First, we trained the separate baseline models on the Logos, AUTSL, and WLASL datasets using the same setup (Section 4). Then, we examined the applicability of the Logos pre-trained model for transfer learning to smaller AUTSL and WLASL datasets. With the model backbone initialized from the Logos pre-train, we evaluated two transfer learning strategies: (a) training all model weights and (b) freezing the pre-trained encoder and training only the classification head. The Logos pre-train substantially improves the model accuracy compared to training from scratch (Table 2).
@@
These experiments demonstrate the possibility of transfer learning from the extensive Logos dataset to other sign languages with only a limited amount of training data.
5.2 Cross-lingual Multi-dataset Co-training
@@
A single model, produced by a multi-dataset co-training, far surpasses the accuracy of the models, separately trained on low-resource datasets from scratch, and also surpasses individual models trained using conventional transfer learning (Table 2). Moreover, results for the WLASL dataset are far above existing SOTA metrics555according to https://paperswithcode.com/ and other papers referring to the datasets in question, see Table 4. As for the AUTSL dataset, note that all leading models use ensembling, pose recognition, and depth maps (or some of the above). In contrast, our model uses a single stream that takes only RGB input.
@@
BSL-1K albanie2020BSL1K 46.9
SignBERT hu2021SignBERT 54.7
SAM-SLR jiang2021SkeletonAwareMultimodal 98.5 58.7
One Model is Not Enough1 96.4
ZBEST zhao2023BESTBERTPretraining 54.6
SignBERT+ hu2023signbert 55.6
NLA-SLR zuo2023NaturalLanguageAssistedSign 61.3
SL-GDN miah2023MultistreamGeneralGraphbased 96.5
ST-GCN2 96.7
Audio-visual … ryumin2023AudiovisualSpeechGesture 98.6
HWGAT patra2024HierarchicalWindowedGraph 95.8 48.5
StepNet shen2024StepNetSpatialtemporalPartaware 61.2
Uni-Sign li2025UniSignUnifiedSign 63.5
MViTv2 (our baseline) 96.58 60.88
Multi-dataset with MM-WLAuslan 97.43 64.59
Multi-dataset with Logos (ours) 97.81 66.82
@@ 5.3 The Encoder Generalization Ability Check
We examined the hypothesis that an encoder pre-trained on the Logos dataset does not produce universal sign features but can only recognize the signs of the pre-train language. When applied to another language, the model maps these signs to the most similar target language signs, as in the approach of wei2023ImprovingContinuousSign. To emulate this hypothesis, we processed the WLASL train set with the Logos pre-trained model and built the map by associating the assigned Logos labels with the most frequent WLASL ground truth labels. Then, we applied the same model to the WLASL test set and substituted the resulting Logos labels with WLASL labels using the map instead of training a target language classification head. We repeated the same experiment with the AUTSL dataset.
@@
5.4 The Importance of the Pre-train Dataset
Pre-train
dataset
Top-1 accuracy
@@ 5.4 The Importance of the Pre-train Dataset
WLASL 93.16 67.90
$M$-WLAuslan 93.64 69.79 45.21 33.01
Table 6: The importance of the pre-train dataset for cross-language transfer learning. Results for both whole and truncated versions of the AUTSL and WLASL datasets using pre-training on the Logos, WLASL, AUTSL, and $M$-WLAuslan datasets.
@@
We also compared the impact of the Logos dataset and the large-scale $M$-WLAuslan dataset on multi-dataset co-training and transfer learning (Tables 4,6). The results obtained with Logos are substantially better than those with MM-WLAuslan for both tasks.
@@ 5.5 The Effect of VSSigns Grouping
VSSigns Top-1 accuracy
train test Whole non-VSSigns VSSigns
Yes Yes 97.90 97.49 98.33
No Yes 97.44 97.10 97.79
No No 87.02 97.10 76.51
@@
Logos
pre-train
on VSSigns
Top-1 accuracy
@@
We investigated the contribution of our approach with grouping labels of visually similar signs in obtaining a high-quality encoder. We trained the classifier on the Logos dataset, using unique pairs of ungrouped and grouped labels as classes. It formed 2,863 ungrouped gloss classes instead of 2,004 grouped VSSign classes in the baseline Logos annotation. Each ungrouped label has a unique associated grouped label, so the model, trained on the ungrouped labels, can be evaluated on grouped labels.
@@
Table 7 shows the accuracy for models trained on 2,004 VSSign and 2,863 ungrouped classes. Quite predictably, the last model yields lower accuracy, primarily due to confusion of VSSigns. However, it is essential that it achieves lower accuracy on grouped VSSigns classes (the classes on which the 1st model was trained). Notably, the degradation is observed even for signs that are not VSSigns whose labels do not differ in all cases.. Furthermore, Table 8 shows that VSSigns grouping results in more effective transfer learning to other sign languages.
6 Conclusion
@@
The paper examines two aspects of the isolated sign language recognition (ISLR) task: cross-language SL model training, including transfer learning, and approaches to handling visually similar signs (VSSigns). To explore these issues, this work presents Logos, a new publicly available Russian Sign Language dataset, the most extensive ISLR dataset by the number of signers and one of the largest available datasets while also the largest RSL dataset in size and vocabulary. It is shown that a model, pre-trained on the Logos dataset can be used as a universal encoder for other language SLR tasks, including few-shot learning. The cross-language transfer learning methods are evaluated, and it is demonstrated that the method of multi-dataset co-training with multiple language-specific classification heads improves SL models for low-resource datasets the most, compared to the conventional ‘‘pre-train and finetune’’ method. The key feature of the Logos dataset is the explicit annotation of visually similar sign groups. With its use, we show that explicitly grouping VSSign labels benefits trained model quality as a video encoder for downstream tasks, such as transfer learning to other sign languages. Based on the proposed contributions, we outperform current state-of-the-art results for the WLASL dataset and get competitive results for the AUTSL dataset, with a single stream model processing solely RGB video.
Limitations
This work is limited by using the MViT baseline architecture and focusing on cross-language transfer learning in ISLR as the downstream task. To generalize the conclusions, further research is needed involving diverse model architectures, low-resource target datasets, and a broader range of downstream tasks, including continuous sign language recognition. We consider this a promising direction for future work.
The Logos dataset reflects the demographics of the participants involved in its collection, resulting in an unbalanced distribution in terms of age and gender. Additionally, the dataset focuses exclusively on RSL, which limits its direct applicability to more diverse settings. Nonetheless, our findings suggest it can still be effectively leveraged in broader sign language recognition tasks.
@@
Legal and ethical aspects were reviewed and approved by our institution’s legal team. All crowdworkers provided informed consent, authorizing the processing and publication of the collected data. Informed consent was provided via a textual form on the crowdsourcing platform. Since all participants were fluent in written Russian, no interpreter translation was required. To save contributors’ privacy, we use anonymized user hash IDs. We do not restrict the participation of signers under 18, provided parental consent was obtained during registration, in compliance with the Civil Code of the Russian Federation. Participation was voluntary. Compensation for completed tasks was aligned with the average salary of a sign language interpreter, proportionate to the time invested. We have verified that the Slovo dataset, incorporated into Logos, adheres to these ethical standards. The dataset is made available exclusively for research purposes. Nonetheless, we acknowledge the potential misuse, such as identifying individuals or enabling large-scale surveillance.
Appendix A Logos Dataset
This appendix contains more detailed characteristics of the Logos dataset and additional technical details for some production steps.
A.1 Dataset Characteristics
Videos have a resolution of at least 720 pixels on the minimum side at a 30 FPS frame rate. About 62% of videos are in FullHD format. Distributions of some dataset characteristics are represented in Figure 3. Among the 381 dataset signers, 41% are 30-40 years old, and 88% are female. We do not limit crowdsourcers by age and gender, and such an uneven distribution reflects the demographics of signers who wish to participate in the project. The participants include deaf individuals, professional interpreters, sign language teachers, and CODA/SODA. Since the data collection was conducted via a crowdsourcing platform, we do not have distribution statistics for these groups. However, all participants passed an RSL proficiency test before contributing. The dataset is divided into 80.7% in the train and 19.3% in the test sets.
Figure 3: Dataset characteristics and distribution analysis. a) Sign length distribution. b) Distance distribution. The distance (in meters) is approximately estimated based on the length between the left and right shoulders of the signer obtained using MediaPipe mediapipe. c) Signers’ devices. d) Devices resolution. e) Number of videos per signer. f) Brightness distribution. The sample brightness is the mean pixel brightness of grayscaled video frames. g) Signers’ gender; h) Signers’ age. The age is determined by the MiVOLO model kuprashevich2023mivolo.
Algorithm 1 Balanced train-test split
#Notation:
$\{G\}:\text{all gloss labels}$
$\{S\}:\text{all signers}$
$\{T\}\subset\{S\}:\text{test set signers}$
$N_{g}:\text{samples number of gloss }g$
$N_{g,s}:\text{ samples number of }g\text{ from signer }s$
$R_{g,s}=N_{g,s}/N_{g}$
#Initialization:
$p\leftarrow 0.2$ $\triangleright$ target test ratio
$\{T\}\leftarrow\text{random }p\text{ from }\{S\}$ $\triangleright$ test set signers
#Optimization:
repeat
  $\forall g:D_{g}\leftarrow\sum_{s\in T}{R_{g,s}}$ $\triangleright$ test gloss ratios
  $d_{wst}\leftarrow\max_{g}{\left|D_{g}-p\right|}$ $\triangleright$ worst deviation
  $g_{wst}\leftarrow{\arg\!\max_{g}}{\left|D_{g}-p\right|}$ $\triangleright$ worst gloss
  #Build sorted list of
  #test signer candidates for replacement:
  if $D_{g_{wst}}>p$ then
   $U\leftarrow\text{sorted}_{\text{desc }}T\text{ by }R_{g_{wst},s\in T}$
  else
   $U\leftarrow\text{sorted}_{\text{asc }}T\text{ by }R_{g_{wst},s\in T}$
  end if
  for $s^{\prime}\text{ in }U$ do
   #Try to replace $s\prime$
   #with signer not in test set:
   for $s^{\prime\prime}\text{ in }S\setminus T$ do
     $\forall g:D^{\prime}_{g}\leftarrow D_{g}-R_{g,s^{\prime}}+R_{g,s^{\prime\prime}}$
     $d^{\prime}_{wst}\leftarrow\max_{g}{\left|D^{\prime}_{g}-p\right|}$
     if $d^{\prime}_{wst}<d_{wst}$ then
      #replace the $s^{\prime}$ signer in $\{T\}$
      $\{T\}\leftarrow\{T\}\cup\{s^{\prime\prime}\}\setminus\{s^{\prime}\}$
      break to outer Repeat
     end if
   end for
  end for
until converge
$\text{\{test video samples\}}\leftarrow\{\text{video: signer}\in T\}$
Output: {test video samples}
A.2 Data Collection
Logos data collection pipeline follows the Slovokapitanov2023SlovoRussianSign pipeline in signer selection, video collection, video validation, and sign time interval annotation stages. The dataset was collected on the crowdsourcing platforms ABC Elementary and Yandex Toloka.
Signers Selection
All project signers confirmed their Russian Sign Language (RSL) proficiency by passing an exam on the crowdsourcing platform ABC Elementary [^6] https://elementary.activebc.ru/ or Yandex Toloka [^7] https://platform.toloka.ai/. The exam involved watching a video demonstrating an RSL sign and selecting the correct translation into Russian. Successful completion required correctly answering at least 17 out of 20 questions.
Video Collection.
Signers watched a video of a correctly performed sign (video template) taken from SpreadTheSign [^8] https://spreadthesign.com/ru.ru/search/ and then recorded a video replicating that sign. Before starting the tasks, signers reviewed the rules for video recording: 1) the gesture must match the example sign; 2) hands must remain within the frame; 3) only one person may appear in the video; 4) video does not shake; 5) video without processing; 6) video must have a short side of at least 720 pixels. Signers could record videos using a smartphone or webcam or upload pre-recorded videos from memory, watch the video, and overwrite the sign.
Video Validation.
At least three validators (SL experts: teachers, professional interpreters) who have successfully completed the validation training and exam will review the video to ensure compliance with the video recording rules. Honeypots (predefined videos with known answers to the customer) were deliberately added to the validation project to identify and eliminate unscrupulous validators.
Time Interval Annotation.
Time interval annotations were necessary to identify the start and the end of the sign dynamics and to exclude unrelated signers’ movements. Each video was pre-processed to a frame rate of 30 fps, enabling the linkage of time intervals to frame numbers. All annotators have completed training and exams on time interval annotation. Honeypots were incorporated into the main tasks to identify and exclude dishonest annotators. Three annotators annotated each video, and the results were aggregated by open-source AggMe framework [^9] https://github.com/ai-forever/aggme.
A.3 Train-test Split
Given that the number of signs recorded by different signers differs, we applied a dynamic programming algorithm to find the approximate solution that has non-intersecting signers in train and test subsets and simultaneously provides approximately 20/80 test/train split ratio both for signers and for each sign samples.
The main idea is that we always select strictly 20% signers and form the test set by the samples they demonstrated. We start with randomly selected 20% signers. Then, in a cycle, we find a sign with a test/train balance the most different from the target 20/80 ratio and try to swap a signer from the test set and a signer not in a test set to reduce the worst disbalance, starting from the test set signers who recorded the largest number of the scoped sign samples. So, we iteratively minimize the maximum deviation from the target test/train sign samples split ratio, keeping the test/train split ratio for signers always equal to the required value. See Algorithm 1 for more details.
Appendix B Sign Language Recognition Pipeline
This appendix provides additional details of the training pipeline used in the work.
B.1 Data Pre-processing
We randomly sample a 32-frame chain from a video with a step of 2 frames. If a sign on a video is longer than 63 frames, the sample is randomly selected within the sign duration. If extra frames are present on the video before and/or after the sign, we extend the range for sample selection up to 5 frames before and after the sign boundaries. For signs shorter than 63 frames, the sample is selected from the sign start and padded at the end by the last frame.
At the training stage Speed Up&Slow Down and Random Add&Random Drop frame sampling augmentations from Kvanchiani et al. pipeline kvanchiani2024TrainingStrategiesIsolated are applied. With probability $p=0.25$, video is accelerated twice or slowed twice with the same $p$. With $p=0.5$ we randomly drop 10% of frames. With $p=0.25$ we truncate a sampled frame chain by 30% and stretch it to the original size by random repeat of remaining frames.
Sampled frames are resized to 300 pix over the longest side and randomly cropped to $224\times 224$ with square padding if needed. Frames are augmented with ColorJitter, RandomNoise, Sharpness, Flip, RandomErasing, and ImageCompression image augmentations. Augmentation parameters are set the same for every frame in a video sample.
Also we use CutMix yun2019CutMixRegularizationStrategy and Mixup zhang2018MixupEmpiricalRisk inter-sample regularization strategies.
B.2 Training Schedule
Training on the Logos dataset was performed on 4 Tesla H100s with 80GB RAM using batch size 16 per GPU for 50 epochs, which took about 40 hours.
For the first 5 epochs, the learning rate linearly increases from 8e-6 to 4.8e-3. Then, a cosine scheduler is used for epochs 6 to 40, reducing LR to 8e-5. Then, LR remains constant.
We use the AdamW optimizer with weight decay=0.05.
When datasets other than Logos are used, or in case of multi-dataset training, to maintain comparable training conditions, including training time, we scale training epochs number and LR schedule to keep the same number of iterations. For instance, for training on a 50% subset of Logos, we train for 100 epochs, with cosine LR annealing from epoch 11 to epoch 80.
For tasks of transfer learning with a frozen encoder, we use a faster training schedule with reduced maximum LR: 15 total epochs with linear LR warm-up from 8e-6 to 8e-4 for 5 times shorter period, 3.5 times shorter cosine annealing from 8e-4 to 8e-5, total training iteration number – 30% of initial training duration. The specified number of epochs is calculated for the target dataset as described above.